# Model-Agnostic Gender Bias Control for Text-to-Image Generation via Sparse Autoencoderthanks: Under Review

### v1

@@ 1 Introduction
Text-to-image (T2I) diffusion models [1, 2, 3, 4] have shown significant progress in generating images with high fidelity and resolution in recent studies. However, potential bias may exist in the generating process regarding sensitive attributes such as gender, race, and age [5, 6, 7]. Among these, gender bias has long been present and remains challenging to eliminate, as it is often implicit and embedded in various semantic contexts. Due to the different social roles traditionally assigned to men and women, gender bias becomes particularly evident in workplace scenarios, and such patterns are learned and reinforced by T2I models [8]. In particular, professional identity prompts frequently expose gender biases in T2I models [9, 10], which reflect and reinforce real-world stereotypes. Controlling such biases during generation remains a challenging problem, particularly without the need for costly model retraining or fine-tuning.
@@
Prior efforts to mitigate gender bias in text-to-image generation can be broadly categorized into training-based and training-free approaches. Training-based methods involve retraining or fine-tuning the model or adding trainable components, which are costly and usually not model-agnostic [11, 12, 13, 14]. Other training-free methods [15, 16, 17, 18] focus on modifying prompts or simply removing gender-related features from the text encoder, but these often lack fine-grained control and may result in the loss of semantic fidelity. Moreover, they tend to overlook the interdependence between features and the inherent ambiguity of semantics, which is particularly problematic when addressing subtle and implicit forms of bias, such as gender bias.
To address these limitations, we propose SAE Debias, a lightweight, model-agnostic framework for mitigating gender bias in profession-related prompts (Figure 1). Our method requires only a one-time training of a sparse autoencoder (SAE) [19] on hidden states from a frozen CLIP-based text encoder. During inference, we identify and subtract gender bias directions in the sparse space—without modifying the original model—thus achieving targeted debiasing with minimal impact on image quality. Our contributions can be summarized as follows:
@@ 2.1 Gender Bias in Stable Diffusion and Debiasing Approaches
Text-to-Image Diffusion Models are inherently biased as they are trained on large datasets that are sampled in real-world settings [20]. Since many T2I models, such as Stable Diffusion (SD) and DALL·E 2 [3], rely on CLIP (Contrastive Language-Image Pre-Training) [21] for text-image alignment, biases in CLIP can propagate into the generated images [22]. [21] found CLIP model exhibits biases against certain protected attributes, and [22] observed that CLIP directs more attention to the physical features of women (such as hair) and exhibits substantial racial and gender bias. Biases in both the text encoder and the training data can propagate into diffusion models, leading to stereotypical outputs during the generation process [23]. As a result, recent research has increasingly focused on identifying and mitigating gender bias in diffusion models.
A common strategy is model retraining or fine-tuning. Esposito et al. [24] fine-tune Stable Diffusion models using debiased synthetic datasets. [25] leverages transfer learning to adapt biased generative models toward fair distributions by fine-tuning on small, unbiased datasets. DeAR [14] extends adversarial training by introducing an additive residual branch to neutralize sensitive attributes. However, such retraining-based approaches are often computationally expensive and lack adaptability to new tasks.
Other methods aim to steer internal representations without retraining. SFID [18] applies low-confidence feature interpolation to trim gender-related features. CLIP-Clip [15] prunes latent features with high mutual information with protected attributes. [16] relies on predefined prompts to mitigate gender bias using a calibrated projection matrix. These methods offer more lightweight solutions, but operations such as direct interpolation and pruning can be overly aggressive, disturbing feature representations and degrading generation quality.
@@ 2.2 Sparse Autoencoders for Interpretability and Bias Control
Sparse Autoencoders (SAEs) [26, 27] have recently emerged as a powerful tool for enhancing interpretability in large models, particularly language and vision-language models. In LLM, SAEs address the inherent polysemanticity problem in neural representations, where a single neuron responds to multiple unrelated concepts [28]. This phenomenon has also been extended to the domain of VLM in [29] recently to fetch visual concepts by selecting the highest activating images in sparse latents of SAEs.
Meanwhile, SAEs offer practical advantages for feature space interventions without retraining the full model. They enable fine-grained feature space level interventions without retraining the whole model. Modifying activations of specific neurons can steer model outputs toward desired attributes or suppress unwanted ones [30]. This makes SAEs particularly suited for debiasing tasks, where bias-inducing components (e.g., gender, race) can be identified and controlled without impacting unrelated features.
@@ 3.1 Sparse Autoencoder
Sparse Autoencoders [31] are unsupervised neural networks consisting of an encoder and a decoder that are optimized jointly using reconstruction error and a sparsity regularization term. The sparsity constraint on the latent activations encourages the network to activate only a small subset of neurons for any given input. These sparsely activated neurons are referred to as sparse latent variables. The sparsity constraint encourages only a few neurons to be activated for a given input so that the most important representations are captured during optimization. The k-sparse autoencoder [32] directly controls sparsity by only keeping the k largest latents and zeroing the rest, providing a fast way to converge in large models.
[…]
@@
[…]
@@
where $W_{\text{dec}}\in\mathbb{R}^{d\times m}$ is the decoder weight. The primary training objective is to minimize the mean squared error (MSE) reconstruction loss between the input $z$ and reconstructed $\hat{z}$, along with an auxiliary loss [19](see supplementary material) that encourages underutilized latents to contribute.
@@ 3.2 Bias in Bios Dataset
We train our k-SAE on the Bias in Bios dataset [34], a widely-used benchmark for evaluating gender bias in natural language processing models. The training split of Bias in Bios consists of 257,478 short biographies, each labeled with one of 28 professional occupations (e.g., physician, nurse, software engineer) and a binary gender attribute inferred from pronouns. The biographies were collected from Common Crawl [35], an open-access repository of large-scale web-scraped data. The dataset is specifically designed to study how occupational classification models exhibit gender biases.
@@ 4.1 Feature Extraction
[…]
@@ 4.2 Training k-sparse autoencoders
[…]
@@ 4.4 Gender Bias Similarity Mapping
[…]
@@ 5.1 Setup
Unlike previous studies on Stable Diffusion debiasing[36] [37] [38] that focus on a limited set of model instances, we evaluate our method across multiple versions, including Stable Diffusion 1.4, 1.5, 2.1, and SDXL [^1] We use stabilityai/stable-diffusion-xl-base-1.0 for SDXL.. For each model, we extract residual features from the text encoder by hooking the output of the second-to-last transformer layer ($L-1$ layer) [^2] For SDXL, we hook the only main text encoder with 1280 dimensions and the auxiliary text encoder remains unchanged. These residual activations serve as the input to train a $k$-sparse autoencoder.
To systematically control and evaluate for gender bias, we utilize the input prompt from [39] to generate images. we construct text prompts following the template: “a photo of $\mathcal{G}$ who works as a[n] $\mathcal{P}$,” where the gender set $\mathcal{G}=\{\text{a man},\text{a woman},\text{a person}\}$ We use 83 professions as in [39]
@@ 5.2 Evaluation Metrics
As discussed in [18], bias in text-to-image models manifests in two major ways: generating the wrong gender for a prompt with a specific gender and generating biased images for a gender-neutral prompt. We evaluate our method under both of these settings. First, we adopt the approach in [18] to detect the gender of generated images using BLIP-2 [40] by asking the binary question: “Does the person look like a male or a female?” and BLIP-2 returns the detected answer. For prompts that explicitly specify a gender, a fair text-to-image generation model should generate images matching the intended gender regardless of the profession. To quantify this, we compute the Mismatch Rate as:
@@
To compare results with [18], we also calculate the Composite Misclassification Rate $M_{RC}$ by averaging the mismatch rates across male and female subgroups, ensuring a balanced evaluation between genders.
@@ 5.3 Results
[…]
Compared to SFID[18] and CLIP-clip[15], our method significantly reduces gender bias without compromising overall performance. Notably, Prompt Debias[16] not only exhibits high mismatch rates on earlier versions but also severely degrades mismatch performance across all categories in SD 2.1 due to overcompensation toward skewness.
@@
SD 1.4 Base 1.45$\pm$0.95 0.24$\pm$0.51 0.84$\pm$0.42 1.74$\pm$0.92 85.18
Ours 0.12$\pm$0.39 0.00$\pm$0.00 0.06$\pm$0.19 0.14$\pm$0.43 83.49
SD 1.5 Base 0.60$\pm$0.63 0.24$\pm$0.51 0.42$\pm$0.29 0.94$\pm$0.65 83.25
Ours 0.60$\pm$0.85 0.12$\pm$0.38 0.36$\pm$0.51 0.66$\pm$0.94 81.93
SD 2.1 Base 1.45$\pm$1.24 0.12$\pm$0.38 0.78$\pm$0.64 1.60$\pm$1.39 83.01
Ours 0.52$\pm$0.84 0.48$\pm$0.84 0.60$\pm$0.63 1.01$\pm$0.98 82.53
SDXL Base 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 93.98
Ours 1.69$\pm$1.52 0.24$\pm$0.51 0.96$\pm$0.86 1.81$\pm$1.63 87.95
@@
SD 1.4 SFID 1.20$\pm$1.27 0.12$\pm$0.38 0.66$\pm$0.60 1.48$\pm$1.34 84.58
Prompt Debias 0.24$\pm$0.51 99.64$\pm$0.58 49.94$\pm$0.19 111.24$\pm$0.94 99.76
CLIP-clip 1.08$\pm$1.33 0.00$\pm$0.00 0.54$\pm$0.66 1.21$\pm$1.48 84.70
Ours 0.12$\pm$0.39 0.00$\pm$0.00 0.06$\pm$0.19 0.14$\pm$0.43 83.49
SD 1.5 SFID 1.41$\pm$0.49 0.00$\pm$0.00 0.70$\pm$0.25 1.57$\pm$0.55 83.68
Prompt Debias 0.52$\pm$0.95 99.6$\pm$0.85 43.84$\pm$15.05 110.57$\pm$2.03 99.70
CLIP-clip 1.08$\pm$1.20 0.12$\pm$0.38 0.60$\pm$0.75 1.15$\pm$1.20 83.37
Ours 0.60$\pm$0.85 0.12$\pm$0.38 0.36$\pm$0.51 0.66$\pm$0.94 81.93
SD 2.1 SFID 1.20$\pm$0.98 0.48$\pm$0.62 0.84$\pm$0.65 1.25$\pm$1.04 83.86
Prompt Debias 8.92$\pm$5.94 15.39$\pm$11.68 11.39$\pm$5.75 17.84$\pm$11.33 76.57
CLIP-clip 0.60$\pm$0.63 0.36$\pm$0.58 0.48$\pm$0.38 0.93$\pm$0.64 84.70
Ours 0.52$\pm$0.84 0.48$\pm$0.84 0.60$\pm$0.63 1.01$\pm$0.98 82.53
@@ 5.5 Interpreting Bias Control and Generation Quality
Diffusion Attentive Attribution Maps We visualize the cross-attention maps by profession of the generated images before and after debiasing. Note that we use the same prompts and generation settings across both conditions to ensure a fair comparison. Specifically, we adopt DAAM (Diffusion Attentive Attribution Maps)[41] to aggregate cross-attention maps across all denoising steps, allowing us to examine the cumulative influence of each token throughout the generation process. In particular, we focus on the attention distributions for the gender-neutral term person and the profession phrases. While debiasing primarily targets gender attributes, it is important that profession-specific information is preserved.
@@
Text-Image Alignment and Visual Quality Evaluation We further evaluate the generated images before and after applying SAE Debias across all model versions using two standard metrics: CLIP Score [42] and Inception Score (IS) [43]. As shown in Table 4, our method achieves comparable CLIP and IS scores before and after debiasing, suggesting that SAE Debias reduces gender bias without significantly affecting image quality or semantic alignment.

### v2

@@ 1 Introduction
Text-to-image (T2I) diffusion models [rombach2022high, podell2023sdxl, ramesh2022hierarchical, mei2024codi] have shown significant progress in generating images with high fidelity and resolution in recent studies. Beyond general text-to-image generation, recent work has also explored fine-grained controllable or instruction-driven editing in both diffusion and autoregressive models [lan2025flux, xu2025scalar, jin2025scar], demonstrating the broader trend of improving semantic controllability at inference time. However, potential bias may exist in the generating process regarding sensitive attributes such as gender, race, and age [luccioni2023stable, anivcin2022bias, wu2023stable]. Among these, gender bias has long been present and remains challenging to eliminate, as it is often implicit and embedded in various semantic contexts. Due to the different social roles traditionally assigned to men and women, gender bias becomes particularly evident in workplace scenarios, and such patterns are learned and reinforced by T2I models [seshadri2023bias]. In particular, professional identity prompts frequently expose gender biases in T2I models [wu2024stable, girrbach2025large], which reflect and reinforce real-world stereotypes. Controlling such biases during generation remains a challenging problem, particularly without the need for costly model retraining or fine-tuning.
@@
Prior efforts to mitigate gender bias in text-to-image generation can be broadly categorized into training-based and training-free approaches. Training-based methods involve retraining or fine-tuning the model or adding trainable components, which are costly and usually not model-agnostic [gandikota2024unified, kim2023stereotyping, orgad2023editing, seth2023dear]. Other training-free methods [wang2021gender, chuang2023debiasing, friedrich2023fair, jung2024unified] focus on modifying prompts or simply removing gender-related features from the text encoder, but these often lack fine-grained control and may result in the loss of semantic fidelity. Moreover, they tend to overlook the interdependence between features and the inherent ambiguity of semantics, which is particularly problematic when addressing subtle and implicit forms of bias, such as gender bias.
To address these limitations, we propose SAE Debias, a lightweight, model-agnostic framework for mitigating gender bias in profession-related prompts (Figure 1). Our method requires only a one-time training of a sparse autoencoder (SAE) [gao2024scaling] on hidden states from a frozen CLIP-based text encoder. During inference, we identify and subtract gender bias directions in the sparse space—without modifying the original model—thus achieving targeted debiasing with minimal impact on image quality. Our contributions can be summarized as follows:
@@ 2.1 Gender Bias in Stable Diffusion and Debiasing Approaches
Text-to-Image Diffusion Models are inherently biased as they are trained on large datasets that are sampled in real-world settings [schuhmann2022laion]. Since many T2I models, such as Stable Diffusion (SD) and DALL·E 2 [ramesh2022hierarchical], rely on CLIP (Contrastive Language-Image Pre-Training) [radford2021learning] for text-image alignment, biases in CLIP can propagate into the generated images [wolfe2022evidence]. [radford2021learning] found CLIP model exhibits biases against certain protected attributes, and [wolfe2022evidence] observed that CLIP directs more attention to the physical features of women (such as hair) and exhibits substantial racial and gender bias. Biases in both the text encoder and the training data can propagate into diffusion models, leading to stereotypical outputs during the generation process [buolamwini2018gender]. As a result, recent research has increasingly focused on identifying and mitigating gender bias in diffusion models.
A common strategy is model retraining or fine-tuning. Esposito et al. [esposito2023mitigating] fine-tune Stable Diffusion models using debiased synthetic datasets. [teo2023fair] leverages transfer learning to adapt biased generative models toward fair distributions by fine-tuning on small, unbiased datasets. DeAR [seth2023dear] extends adversarial training by introducing an additive residual branch to neutralize sensitive attributes. However, such retraining-based approaches are often computationally expensive and lack adaptability to new tasks.
Other methods aim to steer internal representations without retraining. SFID [jung2024unified] applies low-confidence feature interpolation to trim gender-related features. CLIP-Clip [wang2021gender] prunes latent features with high mutual information with protected attributes. [chuang2023debiasing] relies on predefined prompts to mitigate gender bias using a calibrated projection matrix. These methods offer more lightweight solutions, but operations such as direct interpolation and pruning can be overly aggressive, disturbing feature representations and degrading generation quality.
@@ 2.2 Sparse Autoencoders for Interpretability and Bias Control
Sparse Autoencoders (SAEs) [templeton2024scaling, wang2024interpret] have recently emerged as a powerful tool for enhancing interpretability in large models, particularly language and vision-language models. In LLM, SAEs address the inherent polysemanticity problem in neural representations, where a single neuron responds to multiple unrelated concepts [huben2023sparse]. This phenomenon has also been extended to the domain of VLM in [pach2025sparse] recently to fetch visual concepts by selecting the highest activating images in sparse latents of SAEs.
Meanwhile, SAEs offer practical advantages for feature space interventions without retraining the full model. They enable fine-grained feature space level interventions without retraining the whole model. Modifying activations of specific neurons can steer model outputs toward desired attributes or suppress unwanted ones [kim2025concept]. This makes SAEs particularly suited for debiasing tasks, where bias-inducing components (e.g., gender, race) can be identified and controlled without impacting unrelated features.
@@ 3.1 Sparse Autoencoder
Sparse Autoencoders [ng2011sparse] are unsupervised neural networks consisting of an encoder and a decoder that are optimized jointly using reconstruction error and a sparsity regularization term. The sparsity constraint on the latent activations encourages the network to activate only a small subset of neurons for any given input. These sparsely activated neurons are referred to as sparse latent variables. The sparsity constraint encourages only a few neurons to be activated for a given input so that the most important representations are captured during optimization. The k-sparse autoencoder [makhzani2013k] directly controls sparsity by only keeping the k largest latents and zeroing the rest, providing a fast way to converge in large models.
Given the hidden states extracted from the text encoder of the SD model, k-SAE is trained to obtain a sparse latent representation, which captures patterns related to gender and profession in our case. Given the residual features $z\in\mathbb{R}^{n\times d}$ extracted from the (L1)(L-1)-th layer of the text encoder, the k-SAE encoder $f(\cdot)$ maps $z$ into a sparse latent code $h$:
@@
where $W_{\text{enc}}\in\mathbb{R}^{m\times d}$, $b_{\text{enc}}\in\mathbb{R}^{m}$ and are the encoder weights and bias, $b_{\text{pre}}\in\mathbb{R}^{d}$ is actually the decoder bias which is initialized as the geometric median of the dataset according to [bricken2023monosemanticity] to facilitate the recovery of sparse features. $m/d$ is referred to as the expansion factor controlling the hidden layer size relative to the input dimension, and $\text{Top-}k(\cdot)$ retains only the $k$ largest activations [makhzani2013k] and sets the rest to zero.
@@
where $W_{\text{dec}}\in\mathbb{R}^{d\times m}$ is the decoder weight. The primary training objective is to minimize the mean squared error (MSE) reconstruction loss between the input $z$ and reconstructed $\hat{z}$, along with an auxiliary loss [gao2024scaling](see supplementary material) that encourages underutilized latents to contribute.
@@ 3.2 Bias in Bios Dataset
We train our k-SAE on the Bias in Bios dataset [de2019bias], a widely-used benchmark for evaluating gender bias in natural language processing models. The training split of Bias in Bios consists of 257,478 short biographies, each labeled with one of 28 professional occupations (e.g., physician, nurse, software engineer) and a binary gender attribute inferred from pronouns. The biographies were collected from Common Crawl [commoncrawl], an open-access repository of large-scale web-scraped data. The dataset is specifically designed to study how occupational classification models exhibit gender biases.
@@ 4.1 Feature Extraction
where $H=\{h_{1},h_{2},\ldots,h_{n}\}\in\mathbb{R}^{n\times d}$, with $n$ tokens and hidden dimension $d$. We extract the residual features $z$ from the output of the (L1)(L-1)-th layer: The input prompts $X$ are from the Bias in Bios dataset [de2019bias], which consists of textual biographies labeled with binary gender attributes and professional occupations. Specifically, we use the hard text column of each biography as the prompt.
@@ 4.2 Training k-sparse autoencoders
During training, we enforce top-$k$ sparsity on the encoder outputs by retaining only the $k$ largest activations. We experiment with different $k$ values ($k\in\{32,48,64\}$) for different SDs to control the level of sparsity with regard to model size. Specifically, we set the sparsity level to $k=32$ and the expansion factor to $32\times$ for Stable Diffusion 1.x models. For 2.x models, the sparsity level remains $k=32$, but the expansion factor is increased to $48\times$. For SDXL, we adopt $k=64$ with an expansion factor of $64\times$ to accommodate its larger model architecture.
The k-SAE is trained using the combined reconstruction and auxiliary losses defined in Section 3.1. We set the auxiliary loss weight $\alpha$ to $1/32$ as suggested by [kim2025concept] throughout all experiments. During inference, we remove the top-$k$ operation, allowing continuous steering of the latent activations without hard sparsity constraints [kim2025concept], as the Top-K function disregards certain minor activations that may contain important information for both maintaining visual quality and effective debiasing.
@@ 4.4 Gender Bias Similarity Mapping
Here, $i$ refers to each stored profession individually when computing its similarity and weight, while $j$ sums over all stored professions to normalize the softmax. Specifically, $i$ selects the profession whose weight $w_{i}$ is being calculated, and $j$ ensures that the softmax denominator covers all professions for proper normalization. where $T$ is a temperature hyperparameter, and $\text{Sim}(\cdot,\cdot)$ denotes cosine similarity.
@@ 5.1 Setup
Unlike previous studies on Stable Diffusion debiasing[jiang2024mitigating] [he2024debiasing] [huang2025debiasing] that focus on a limited set of model instances, we evaluate our method across multiple versions, including Stable Diffusion 1.4, 1.5, 2.1, and SDXL [^1] We use stabilityai/stable-diffusion-xl-base-1.0 for SDXL.. For each model, we extract residual features from the text encoder by hooking the output of the second-to-last transformer layer (L1L-1 layer) [^2] For SDXL, we hook the only main text encoder with 1280 dimensions and the auxiliary text encoder remains unchanged. These residual activations serve as the input to train a $k$-sparse autoencoder.
To systematically control and evaluate for gender bias, we utilize the input prompt from [cho2023dall] to generate images. we construct text prompts following the template: “a photo of $\mathcal{G}$ who works as a[n] $\mathcal{P}$,” where the gender set $\mathcal{G}=\{\text{a man},\text{a woman},\text{a person}\}$ . We use 83 professions as in [cho2023dall]
@@ 5.2 Evaluation Metrics
As discussed in [jung2024unified], bias in text-to-image models manifests in two major ways: generating the wrong gender for a prompt with a specific gender and generating biased images for a gender-neutral prompt. We evaluate our method under both of these settings. First, we adopt the approach in [jung2024unified] to detect the gender of generated images using BLIP-2 [li2023blip] by asking the binary question: “Does the person look like a male or a female?” and BLIP-2 returns the detected answer. For prompts that explicitly specify a gender, a fair text-to-image generation model should generate images matching the intended gender regardless of the profession. To quantify this, we compute the Mismatch Rate as:
@@
To compare results with [jung2024unified], we also calculate the Composite Misclassification Rate $M_{RC}$ by averaging the mismatch rates across male and female subgroups, ensuring a balanced evaluation between genders.
@@ 5.3 Results
As shown in Table 2, our method consistently outperforms other approaches across all metrics on SD 1.4 and achieves the best skew values across all models. It obtains either the lowest or the second-lowest mismatch rates for both male and female prompts, reaching an overall mismatch rate of just $0.06\%$ on SD 1.4 while maintaining strong performance on SD 1.5 and SD 2.1.
Compared to SFID[jung2024unified] and CLIP-clip[wang2021gender], our method significantly reduces gender bias without compromising overall performance. Notably, Prompt Debias[chuang2023debiasing] not only exhibits high mismatch rates on earlier versions but also severely degrades mismatch performance across all categories in SD 2.1 due to overcompensation toward skewness.
@@
[…]
SDXL Base $\pm0.00$ $\pm0.00$ $\pm0.00$ $\pm0.00$ 93.98
[…]
@@
[…]
@@ 5.5 Interpreting Bias Control and Generation Quality
Diffusion Attentive Attribution Maps We visualize the cross-attention maps by profession of the generated images before and after debiasing. Note that we use the same prompts and generation settings across both conditions to ensure a fair comparison. Specifically, we adopt DAAM (Diffusion Attentive Attribution Maps)[tang2022daam] to aggregate cross-attention maps across all denoising steps, allowing us to examine the cumulative influence of each token throughout the generation process. In particular, we focus on the attention distributions for the gender-neutral term person and the profession phrases. While debiasing primarily targets gender attributes, it is important that profession-specific information is preserved.
@@
Text-Image Alignment and Visual Quality Evaluation We further evaluate the generated images before and after applying SAE Debias across all model versions using two standard metrics: CLIP Score [hessel2021clipscore] and Inception Score (IS) [salimans2016improved]. As shown in Table 4, our method achieves comparable CLIP and IS scores before and after debiasing, suggesting that SAE Debias reduces gender bias without significantly affecting image quality or semantic alignment.